OpenStudy (anonymous):

Solve x^2 - 3x = -8.

OpenStudy (anonymous):

Solve this quadratic equation like this : First try to convert it in form ax^2+bx+c=0 where a,b,c are parameters(constant) and x is variable So we do in this question take 8 to right hand side we get x^2-3x+8=0 Take the D the discriminant= (b^2-4*a*c)^1/2 =(-3^2-4*1*8)^1/2 =-23^1/2= \[\sqrt{-23}=i*\sqrt{23}\] i is the value =\[\sqrt{-1}\] Now we have 2 roots as it is quadratic as r1=(-b+D)/4a and r2=(-b-D)/4a so r1=(3+i* \[\sqrt{23}\])/4 and r2=(3-i* \[\sqrt{23}\])/4